A total of
724
tickets were sold for the school play. They were either adult tickets or student tickets. There were
74
more student tickets sold than adult tickets. How many adult tickets were sold?

Answer by macston(5194)

A=number Adult tickets;S=A+74= number Student tickets
Total tickets=adult tickets + student tickets
724=A+(A+74)
724=2A+74 Subtract 74 from each side
650=2A divide each side by 2
325=A ANSWER 1: There were 325 Adult tickets sold
S=A+74=325+74=399 ANSWER 2: There were 399 student tickets sold